OpenStudy (anonymous):

if log(M/N)=4 and log(P/N)=5, what can you say about the relationship between M and P?

OpenStudy (anonymous):

hint: log(M/N)=4 and log(P/N)=5 then 5 log(M/N) = 4log(P/N) log ((M/N) ^5) = log((P/N)^4) ((M/N) ^5) = ((P/N)^4) and so on...